OpenStudy (anonymous):

The ages of a large group of people have an average of 52 years and an SD of 6 years. The average annual income of the people is $48,000 and the SD is $27,000. The correlation between the two variables is 0.45. 1A This problem is worth 1 point. The r.m.s. error of regression for estimating income based on age is closest to: 5.4 thousand dollars 5.4 years 24.11 thousand dollars 24.11 years

OpenStudy (anonymous):

did any one had solve it

OpenStudy (kropot72):

The r.m.s error of the regression line is \[\sqrt{(1-r ^{2}})\times S.D.\ of\ income=\sqrt{(1-0.45^{2}}\times 27000=$24,111\]

OpenStudy (anonymous):

thank you very much

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

If you have any questions allso feel free bro

OpenStudy (anonymous):

Are you also member of edx

OpenStudy (kropot72):

No, I am not a member of edx.

OpenStudy (anonymous):

Is the scatter diagram of the two variables football shaped? No. Yes. Maybe, or maybe not; there is not enough information to decide.

OpenStudy (kropot72):

A scatter diagram is football-shaped or the points form a cloud on the scatter diagram if most of the points on the graph fall within the region of an elliptical shape. The annual income has a very large standard deviation relative to the average, Therefore the scatter diagram is not football-shaped.

OpenStudy (anonymous):

Thanks very good exlonation

OpenStudy (anonymous):

One question sir you are from where

OpenStudy (anonymous):

The weights of newborns at a hospital have an average of 7 lb and an SD of 1.2 lb. The average head circumference of the newborns is 14 inches and the SD is 0.5 inches. The correlation between the two variables is 0.55 and the scatter diagram is football shaped. The regression line for estimating head circumference based on weight is drawn on the scatter diagram, with weight on the horizontal axis. 3A : 2.0 POINTS For about 50% of the newborns, the regression estimate of head circumference based on weight is correct to within _____________ inches. Fill in the blank.

OpenStudy (kropot72):

If the scatter diagram is football-shaped, the r.m.s. error can be used like a standard deviation for the regression line. Approximately 50% of the points will be between \[the\ regression\ estimate \pm 0.675(r.m.s.\ error)\] \[r.m.s.\ error=\sqrt{(1-r ^{2})}\times S.D.\ head\ circumference=0.835\times 0.5=0.4175\ inch\] For about 50% of the newborns, the regression estimate of head circumference based on weight is correct to within (0.675 * 0.4175) inch.

OpenStudy (anonymous):

For about what percent of the newborns does the regression line overestimate head circumference by more than 0.5 inches? 3C : 3.0 POINTS Of the newborns who weigh 8 pounds, about what percent have head circumferences larger than 15 inches?

OpenStudy (kropot72):

For about what percent of the newborns does the regression line overestimate head circumference by more than 0.5 inches? z * (r.m.s. error) = 0.5 z = 0.5/0.4175 = 1.1976 Reference to a standard normal distribution table shows that about 11.5% will have head diameters overestimated.

OpenStudy (anonymous):

answears are wrong bro

OpenStudy (anonymous):

sorry they are right what about 15 inches

OpenStudy (anonymous):

what about part c

OpenStudy (anonymous):

Had done it already very simple

OpenStudy (anonymous):

The ages of a large group of people have an average of 52 years and an SD of 6 years. The average annual income of the people is $48,000 and the SD is $27,000. The correlation between the two variables is 0.45. 1A This problem is worth 1 point. The r.m.s. error of regression for estimating income based on age is closest to: 24.11 thousand dollars(wright answear) 1B : 2.0 POINTS Is the scatter diagram of the two variables football shaped? Yes.(right answear) Problem 2A The calculations are wrong.(right answear) Problem 2B There is a non-linear relation between the two variables in the regression. (right answear) Problem2C Regresion looks ok (right answear) Problem 2D The regression looks OK(right answear) Problem 2E The calculations are wrong.(right answear) 3A : 2.0 POINTS For about 50% of the newborns, the regression estimate of head circumference based on weight is correct to within _____________ inches. Fill in the blank. 0.28181 correct 3B : 2.0 POINTS For about what percent of the newborns does the regression line overestimate head circumference by more than 0.5 inches? 11.5 correct 3C : 3.0 POINTS Of the newborns who weigh 8 pounds, about what percent have head circumferences larger than 15 inches? 3.2 correct

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

An orchestra consists of 25 members. The youngest member is 32 years old. She leaves and is replaced by a new member who is 30 years old. 1A By how much does the replacement change the average age of the members of the orchestra? [Note on answer format: Calculate your answer in years. If you think the average is unchanged, enter 0. If you think it goes up by 10.5 years, enter 10.5. If you think it goes down by 10.5 years, enter -10.5.] is -10.5???

OpenStudy (anonymous):

@rocio.soto estas son las respuestas: A -0.08 B 0

OpenStudy (anonymous):

ThanK You

OpenStudy (anonymous):

Please, Help me with this questions The final exam is “graded on a curve,” with the top 20% of the students on the final receiving an A grade on the exam. Is the set of students who receive an A on the exam based on the new final exam scores different from the set of students who received an A on the exam based on the old final exam scores? The correlation between the new final scores and the old final scores is: -1 -0.63 0 0.63 1 impossible to determine based on the information given. possible to determine based on the information given, and not equal to -1, -0.63, 0, 0.63 or 1. The correlation between the new final scores and the midterm scores is: -1 -0.63 0 0.63 1 impossible to determine based on the information given. possible to determine based on the information given, and not equal to -1, -0.63, 0, 0.63 or 1.

OpenStudy (anonymous):

the answer is 1 and next is 0.63. please help me with this questionThe scatter diagram of midterm scores and final scores in a large class is football shaped. For about 80% of the students, the regression estimate of final score based on midterm score is correct to within 15 points. For about 95% of the students, the regression estimate of final score based on midterm score is correct to within ___________ points.

OpenStudy (anonymous):

A standard deck consists of 52 cards, distributed as follows: There are 13 cards in each of 4 suits: hearts, diamonds, spades, and clubs. Hearts and diamonds are red. Spades and clubs are black. In each suit there are 13 ranks: Ace, 2, 3, 4, 5, 6, 7, 8, 9, 10, Jack, Queen, King. In Exercises 2A through 2E, assume that cards are dealt from a well-shuffled deck; that is, cards are dealt at random without replacement. PROBLEM 2A This Problem is worth 1 Point The "face" cards are those that have a person’s face on them: Jack, Queen, or King. One card is dealt. Find the chance that it is a face card. PROBLEM 2B This Problem is worth 1 Point One card is dealt. Find the chance that its color is black and its rank is an even number. PROBLEM 2C This Problem is worth 1 Point One card is dealt. Find the chance that its color is red or its rank is an even number. PROBLEM 2D This Problem is worth 1 Point A poker hand (5 cards) is dealt. Find the chance that the cards are all hearts. PROBLEM 2E This Problem is worth 1 Point A poker hand (5 cards) is dealt. Find the chance that the cards are all of the same suit. [This is sometimes called a “flush.”] There are 5 people in a room. Make the following simplifying assumptions about their birthdays: Ignore leap years; assume 365 days in the year. For each person, assume that the chance he/she is born on a specified day is 1/365. No twins! Assume that the chance that someone is born on a specified day is not affected by other people’s birthdays. PROBLEM 4B This Problem is worth 1 Point Find the chance of a "match": that is, at least two people in the room have the same birthday. PROBLEM 4C This Problem is worth 1 Point Repeat 4B for a room in which there are 400 people. There are 26 letters in the English alphabet. Letters are picked one by one at random, so that each letter has the same chance of appearing as any other letter, regardless of which letters have appeared or not appeared. PROBLEM 5A This Problem is worth 1 Point Six letters are picked. Find the chance that the sequence that appears is RANDOM, in that order. PROBLEM 5B This Problem is worth 1 Point Six letters are picked. Find the chance that they can be arranged to form the word RANDOM. A two-day conference is going to be held in a city. The leading newspaper in the city says there is a 10% chance of rain on the first day of the conference and a 40% chance of rain on the second day. PROBLEM 6A This Problem is worth 1 Point If the newspaper’s figures are correct, the chance that it rains in the city during the conference is: at least 0%, but we can’t get a better lower bound with the information given. at least 10% but it might be less than 40%. at least 40%. PROBLEM 6B This Problem is worth 1 Point If the newspaper’s figures are correct, the chance that it rains in the city during the conference is at most 10%. at most 40%. at most 50%. at most 100%; that’s the best upper bound we can get based on the information given.

OpenStudy (anonymous):

2A : 3/13 ???